# Designing Adaptive User Interfaces for mHealth applications targeting chronic disease: A User-Centric Approach

### v1

@@ 3.4.2. Quantitative data analysis
We exclusively gathered quantitative data through our survey, and subsequently performed statistical analysis using Microsoft Excel. Descriptive statistics provide a comprehensive overview of the characteristics and preferences of the survey respondents, allowing us to understand the distribution and trends within the dataset. Chi-square tests were employed to determine if there was a significant association between the preferences of users with respect to various aspects of adaptations and their demographic characteristics, such as age, gender, nationality, education, and chronic diseases. To ensure that the data distribution meets the prerequisites for the Chi-square independence test, related variables were grouped into categorical variables beforehand. Age was categorised into two groups: 18-45 and 45-74. The level of education was classified into three groups: Less than Bachelor’s degree, Bachelor’s degree, and Postgraduate (Master’s and Doctoral degree). Chronic disease conditions were categorised as detailed in Section 5.1. If a significant association is found, binary logistic regression or multinomial logistic regression will be subsequently employed to model the relationship between these variables. Understanding how these demographic factors influence user preferences provides valuable insights into tailoring AUIs to accommodate diverse user needs and preferences. The common significance level of is chosen for statistical analysis.
4. Findings from interviews and focus groups
We present our findings on user’s perspectives towards the AUI prototype. We identified four overarching challenges that participants faced while interacting with our AUI mHealth prototype (outlined in Table 2) . We also found a set of strategies applied or purposed by participants to have a better user experience of the adaptations (outlined in Table 3). These strategies were founded to be moderating by different characteristics of users.
| # | Age | Gender | Ed* | Nationality | Chronic diseases | mHealth app usage pattern |
| Iteration 1: Each focus group study took 60-70 minutes. | ||||||
| FP1 | 55-64 | Female | M | Australian | Type 2 diabetes | Less than once a week/1-10 minutes |
| FP2 | 55-64 | Female | M | Australian | Type 2 diabetes | Less than once a week/20-30 minutes |
| FP3 | 65-74 | Female | M | Australian | Type 2 diabetes | Once a week/10-20 minutes |
| FP4 | 65-74 | Male | P | Australian | Type 2 diabetes | 2 or more times per day/1-10 minutes |
| FP5 | 25-34 | Female | M | Chinese | Asthma, Allergic rhinitis & Chronic gastritis | Less than once a week/1-10 minutes |
| FP6 | 65-74 | Female | P | Tamil | Type 2 diabetes | Less than once a week/1-10 minutes |
| FP7 | 55-64 | Female | T | Australian | Type 2 diabetes | Less than once a week/1-10 minutes |
| FP8 | 18-24 | Female | B | Malaysia | Depression | Less than once a week/1-10 minutes |
| FP9 | 65-74 | Male | B | Australian | Type 2 diabetes, Heart disease, Epilepsy&Asthma | 2 or more times per day/1-10 minutes |
| Iteration 2: Each focus group study took 60-70 minutes. | ||||||
| FP10 | 25-34 | Male | M | Chinese | Type 2 diabetes | 2 or more times per day/1-10 minutes |
| FP11 | 25-34 | Female | B | Filipino | Over weight, Asthma & Anxiety | 2 or more times per day/1-10 minutes |
| FP12 | 25-34 | Male | M | Chinese | Rhinitis & Anxiety | 2 or more times per day/1-10 minutes |
| FP13 | 35-44 | Male | B | India | Overweight & Type 2 diabetes | less than once a week/1-10 minutes |
| FP14 | 35-44 | Male | B | India | Hypertension | less than once a week/1-10 minutes |
| FP15 | 25-34 | Female | M | Chinese | Mental health conditions | 2 or more times per day/more than 30 minutes |
| Iteration 3: Each interview study took 30-60 minutes (aimed at saturation). | ||||||
| IP1 | 55-64 | Female | A | Australian | Type 2 diabetes | 2 or more times per day/1-10 minutes |
| IP2 | 35-44 | Male | B | British | Mental health conditions | Once a week/1-10 minutes |
| IP3 | 65-74 | Female | A | Australian | Type 2 diabetes | 2 or more times per day/1-10 minutes |
| IP4 | 45-54 | Female | M | Australian | Hypertension& Type 2 diabetes | 2 or more times per day/1-10 minutes |
| IP5 | 45-54 | Female | M | Chinese | Anxiety, Insomnia & Mental health conditions | 2 or more times per day/1-10 minutes |
| IP6 | 18-24 | Female | B | American | Anxiety & Mental health conditions | 2 or more times per day/1-10 minutes |
| IP7 | 55-64 | Female | B | Australian | Major chronic depressive disorder & PTSD | 2 or more times per day/1-10 minutes |
| Ed*: Education ( M=Master, B=Bachelor, A=Associate, T=Technical training ) | ||||||
4.1. Focus group and interview participants
Table 1 provides a breakdown of the participant characteristics, showcasing a variety of nationalities such as Chinese, Indian, British, Filipino, Malaysian, American, and Tamil, with a notable majority being Australian (n=10). The sample comprises a greater proportion of females (n=15) than males (n=7), spanning across various age groups (18-34: 7, 35-54: 6, 55-74: 10), as well as a wide spectrum of chronic diseases, such as type 2 diabetes (n=12), rhinitis (n=2), chronic gastritis (n=1), epilepsy (n=1), hypertension (n=2), asthma (n=3), chronic depressive disorder (n=1) and different mental health conditions (n=7). All participants have current or previous experience with mHealth applications. Their experience is assessed based on the frequency and duration of their usage (Robbins et al., 2017). Participants were asked to indicate how frequently they use health apps and to specify the duration of each session. The result indicates varying usage patterns among participants. 8 users reported using the app less than once a week, while 12 mentioned using it two or more times a day. Most users (n=19) reported a session duration of 1-10 minutes. Most of the participants (n=17) hold a university degree. Of these, almost half have obtained a bachelor’s degree (n=8), and the remaining half have earned a master’s degree (n=9). Participants are numbered according to their study involvement: FP1-FP15 for focus group participants and IP1-IP7 for interview participants.
4.2. Challenges participants face in using AUIs
Drawing on STGT data analysis and the existing literature (Wang et al., 2023b; Abrahão et al., 2021), we grouped the identified challenges into four distinct categories underpinned by the key concepts (see Table 2). These categories are as follows: (1) What to adapt: which revolves around identifying the specific elements or aspects of the UI that require adaptation. It involves understanding the potential challenges that may arise when different types of adaptations are implemented; (2) Who should initiate adaptations: This category addresses the roles and responsibilities associated with initiating interface adaptations; (3) How to adapt: This category focuses on the mechanisms and strategies utilised to effectively implement interface adaptations; (4) When to adapt: This category pertains to determining the optimal timing and conditions under which adaptations should take place.
| Categories | Concepts | Participants | # | Moderating effect(NA:Not Applicable) |
| C1. What to adapt (10 FP, 7 IP) | Predictability of the user interface | FP 2, IP 4,5,7 | 4/22 | NA |
| Comprehensibility of adaptation | FP 2,4,7,9,12,15,16, IP 3 | 8/22 | Technological proficiency | |
| Obtrusiveness of generated adaptations | FP 4,5,6,7,9, IP 1,3,4,6,7 | 10/22 | mHealth usage experience | |
| Mismatch of generated adaptations | FP 4, IP 2,3,6 | 4/22 | Self-sufficiency and independence | |
| Aversion to negative measurements | FP 8,9, IP 1-5, | 7/22 | Health conditions&Coping mechanisms | |
| C2. Who should initiate adaptation (12 FP, 6 IP) | System is unable to initate the adaptation | FP 4-6,9,11,12,15, IP 3,6 | 9/22 | NA |
| User might not be the right person to handle the adaptation | FP 1-4,6,7-9,11,12, IP 1-4,6,7 | 16/22 | Cognitive load | |
| C3. How to adapt (4 FP, 3 IP) | Infringement of privacy | FP 11,13, IP 2 | 3/22 | NA |
| Decision fatigue and frustration | FP 7,9, IP 2-4 | 5/22 | NA | |
| C4. When to adapt (6 FP, 6 IP) | Frequency of adaptation | FP 4,6,8,9, IP4,7 | 6/22 | Characteristic of features |
| Timing of adaptation | FP 4,6,7,8,10, IP 1,3,4,6 | 9/22 | Health conditions&mHealth usage experience | |
| Usage scenario | FP 8,9,10, IP 1,3,4,5,7 | 8/22 | NA |
4.2.1. C1. What to adapt
The importance of UI predictability emerged as a significant concern, consistent with prior research findings (Mezhoudi et al., 2015; Jameson, 2007). The participants emphasised the importance of spatial stability within the UI (IP 4,5,7, FP 2). Participants rely on familiar patterns and locations within the UI to locate information and perform tasks efficiently. IP7 articulated, ”system makes lots of changes would confuse me because I get used to finding things in the same place. When it changes, it is hard to see what will come next”. Studies have shown that even minor changes in interface layout or functionality can disrupt users’ mental models and increase the likelihood of errors (Hammond, 1987). Therefore, ensuring stability and predictability in UI design is essential for supporting users’ cognitive processes.
Some participants faced difficulties in understanding the concepts of UI adaptation and the resulting adaptations, despite the assistance and explanations provided by the interviewer or facilitator (FP 2,4,7,9,12,15,16, IP 3). This lack of comprehensibility could potentially hinder the effectiveness of adaptations, leading to a decrease in user performance and satisfaction, and ultimately causing users to consider switching to alternative applications (FP 2). Within the context of chronic diseases, the intricacies of adaptation may be further exacerbated (Yuan and Herbert, 2012). For example, in our AUI prototype, an avatar displayed a concerned expression in response to unfavourable measurements. However, older participants (over 65 years) (FP 4, IP 3) suggested that this simple visual cue may not alleviate their concerns and may even increase their confusion. On the contrary, younger participants (aged less than 44 years) appreciated such avatars, viewing them as a means to ”make all serious things more fun” (IP 7, FP 5, 8, 10, 13). This divergence is consistent with previous research indicating that gamified design is more appealing to younger audiences due to their technological proficiency and familiarity with gaming (Betts et al., 2019). The complexity inherent in the management of chronic diseases requires adaptation strategies that are easily understandable for users of different age groups and backgrounds. Older adults may require straightforward and explicit communication of adaptations. On the contrary, younger users may appreciate more interactive and engaging elements.
Some participants expressed concerns about specific adaptations, particularly those that involve comfort words and gamified design elements to convey health-related information, citing them as obtrusive and potentially diverting attention from primary tasks (IP 1,3,4,6,7, FP 4,5,6,7,9). For example, instead of simply stating the health measurement result in a clinical way, the app could say something like: ”I’m sorry to hear about your recent health measurement result. Remember, setbacks are a normal part of the journey. Let us work together to find ways to improve your health”. Additionally, it is important to consider the broader challenges associated with gamification in mHealth applications, including issues related to gamification appropriateness and socio-cultural adaptability (Vajawat et al., 2021). Gamification affects patients differently, and mismatched adaptations may have a negative influence on patients (Codish and Ravid, 2014). Participants with more experience using mHealth applications tended to view certain distracting adaptations less negatively (FP 5, IP 6). This suggests that familiarity with the mHealth tools may influence how users perceive and interact with various design elements (McDougall et al., 2000).
Mismatched adaptations were also mentioned a lot by the participants (FP 4, IP 2,3,6). They worried about being trapped in an adaptation process triggered by ”wrong measurements or incorrect input data” (FP 4, IP 6). IP2 shared similar feelings, stating that he ”felt overwhelmed due to the uncertainty of what the system would understand and what it would not”. Such instances of expectation mismatches often result in a decrease in trust in the adaptation generated (Glass et al., 2008). Integration of sensor networks and data in mHealth has the potential to improve healthcare services for patients (Albahri et al., 2018). However, instances of ”sensor failure” (IP 3) or ”inaccurate input data” (IP 6) from patients can raise concerns about the generation of unwanted adaptations. Furthermore, some younger participants (aged under 34) acknowledged the significance of certain adaptations, such as adjustments in font size or simpler UI designs (FP 8,11,15, IP 6). However, they often perceived these adaptations as beneficial not for themselves but for individuals with varying disabilities and levels of independence. Their hesitation to fully embrace these adaptations may be attributed to their strong sense of self-sufficiency and independence (Trajkova and Martin-Hammond, 2020), which leads them to overlook the immediate advantages of such adaptations. These perspectives highlight the importance of focusing on long-term benefits or positive impact on individuals, broadening the acceptance of adaptations.
Participants in our study exhibited mixed feelings about the measurement results, particularly in the context of recording health data such as blood glucose levels, stress levels, and blood pressure. Some participants expressed a strong aversion towards bad measurement results and preferred not to be constantly reminded of negative outcomes (IP 1-5, FP 8,9). For example, IP1 shared her tendency to avoid doctors and attempted to deny the measurement when faced with unfavourable health data. Participants suggested that presenting the data in a ”noncritical” (IP 4) and ”positive reinforcement” (IP 3) manner would be beneficial. These struggles highlight the influence of user symptoms and health conditions on their interactions with technologies (Kornfield et al., 2020). However, other participants, particularly those with mental health problems, found value in reminders and clear presentations of such data as important signals for their health management (IP 5, FP 5,8,10-13,15). For example, IP5, who deals with mental health issues, expressed that:
The different attitudes towards measurement results observed among participants suggest different coping mechanisms employed to manage stressful situations (Csikszentmihalyi et al., 2014). Some individuals may adopt avoidance and denial as coping strategies, others may employ more positive coping strategies to reduce stress and help the mind and body return to their original state of functioning (Folkman and Lazarus, 1980).
4.2.2. C2. Who should initiate adaptation
In typical app usage scenarios, users already face the challenge of continuously recording various types of data, such as heart rate, blood pressure, blood sugar levels, and body composition (Philip et al., 2022). The introduction of AUIs in these applications can further complicate matters, primarily due to the crucial role of data in driving UI adaptation (Alvarez-Cortes et al., 2007). Many participants in our study already struggle to maintain complete and accurate daily data records, resulting in occurrences of incomplete data that impede the system’s ability to initate the adaptation process (FP 4-6,9,11,12,15, IP 3,6). For example, FP9 revealed his frequent lapses in recording blood tests, while FP5 acknowledged that he forgot to record her daily mood despite setting up multiple reminders. Users may struggle to enter their health data for various reasons, including forgetfulness, lack of motivation, or difficulty integrating the use of the app into their daily routines (Stawarz et al., 2014). This inconsistency in data recording is a common issue in mHealth applications related to chronic diseases (Short et al., 2018). This problem is often attributed to the low levels of user adherence and the high rates of attrition commonly observed in mHealth interventions (Kelders et al., 2012). Furthermore, IP3 admitted that sometimes ”fakes some data” to complete records artificially, which can ”make the progress and reports look really good”. These participants emphasised that having various adaptations becomes pointless if the underlying data is unreliable. As a result, the reliability and completeness of health records can be compromised, limiting the adaptation of these functions in monitoring and managing chronic disease functions.
Based on the discussion of data collection, participants expressed concerns about whether user is the right person to handle the adaptation process. Firstly, older participants (over 45 years) mentioned the added mental workload of overseeing adaptation processes along with their existing responsibilities related to managing their chronic conditions (IP 1,4,7, FP 4,6,7,9). This additional cognitive load could potentially overwhelm them, hindering their ability to effectively self-manage their health. IP1 noted that:
Added to this, participants also found it challenging to learn and navigate the adaptation process, which could counteract the time and effort savings benefits that automation of adaptations aims to offer (Gajos et al., 2006; Harman et al., 2014). Some participants expressed reservations about complete automation, fearing that it could lead to a lack of awareness about system changes (IP 3,6, FP 1,2,3). This suggests that mHealth apps could pose risks to user safety when their design is not well controlled (Akbar et al., 2020).
Secondly, considering the intended purpose of the app, situations can arise where a user’s preferences conflict with the intended use of the app (FP 4,6,7-9,11,12, IP 2). For example, in the context of rearranging elements on the UI, FP4 expressed concerns that:
FP11 and FP12 suggested that some users might manipulate the system to obtain the adaptations they want, deviating from the original design and purpose of an app. For instance, a user who always forgets to record or report lower activity levels than the actual, hoping to receive suggestions for lighter workouts or fewer reminders to be active. In these cases, users manipulate the system to obtain adaptations that deviate from promoting genuine physical activity and supporting users’ health goals.
4.2.3. C3. How to adapt
mHealth devices are capable of monitoring various aspects, including activities, sleep patterns, emotional states, vital signs such as blood pressure, and mental health conditions (Prasad et al., 2014). Several younger participants (under 44 years of age) in our study raised concerns about potential privacy infringements related to adaptation (IP 2, FP 11,13). IP2 emphasised that he ”probably would not allow any audio-based input for every app” because he was worried that the app is listening all the time. This observation implies that smartphone users may feel violated when discovering that applications access their data without their knowledge (Li et al., 2020), potentially leading individuals with greater privacy concerns to avoid using such apps (Lohr, 2015). On the contrary, other participants expressed confidence in the reputation of the application (FP 1,2,3,4,6) or the organisation (FP 14,15, IP 1,3,5,7) that developed the application, suggesting that they are not overly concerned about privacy issues. This sentiment is deemed reassuring given the prevalent privacy issues surrounding mHealth apps (Haggag et al., 2022). Furthermore, participants stressed the importance of implementing stringent data protection regulations and accountability mechanisms to prevent unauthorised use or disclosure of user data (FP 14,15, IP 2).
Excessive control over adaptation can result in decision fatigue and frustration, especially concerning individuals with chronic diseases already facing making numerous decisions related to their treatment and lifestyle adjustments (Scherer et al., 2021). Requesting participants to decide on adaptations might make them feel more drained as the day progresses (IP 2,3,4, FP 7,9), as this forces them to make more decisions (Sara, 2021). Stress associated with making frequent adaptation decisions can exacerbate existing health conditions and negatively impact overall well-being (Scherer et al., 2021). Participants indicated a preference for a ”system-driven approach” to adaptation decisions (IP 2).
4.2.4. C4. When to adapt
While the participants appreciated the ability to adapt the app, they found it challenging to engage with frequent adaptations (FP 4,6,8,9, IP 4,7), especially if they needed to perform these operations daily. FP6 expressed concerns about the time and effort required to manage constant changes, stating that it became overwhelming to keep up with the app’s adaptations alongside other daily responsibilities. Moreover, certain participants wished the UI to remain unchanged after the initial setup, particularly for less frequently used features (FP 4,6). They acknowledged the necessity of dynamic adjustments for frequently accessed features.
The timing of adaptation is a crucial aspect of UI adaptation (Bunt et al., 2004; Abrahão et al., 2021), and participants in our study expressed varying preferences in this regard. Some participants who were knowledgeable about their health conditions and app usage preferred setting the adaptation after logging in (IP 6, FP 8,10). This aligns with the notion that upfront customisation is efficient when users have a clear understanding of their needs (Bunt et al., 2004). On the contrary, the majority of participants preferred adaptations to occur at regular intervals but should wait until they had familiarised themselves with the application (FP 4,6,7, IP 1,3,4). This suggests that some users prefer a gradual introduction of adaptations to allow them to acclimate to the app’s interface and functionality before making any significant changes. FP7 expanded on this point, emphasising that immediate adaptation upon initial login could be overwhelming, particularly for those who had just been diagnosed with chronic diseases. Understanding user decision-making strategies is crucial, as they are influenced by their level of expertise in a given domain (Bettman, 1971). Expert users, who possess a deeper understanding of the app’s functionalities, can leverage their domain knowledge to exercise more precise control over the system. This highlights the importance of providing support to novice users while offering flexible customisation options for those with greater expertise. Such approaches can empower users of all levels to effectively navigate and leverage the adaptive capabilities of the app (Attfield et al., 2011).
Being attentive to various usage scenarios is essential to ensure that the application remains relevant and beneficial to a diverse user base (Stephan et al., 2017; Varshney, 2014). Understanding how users interact with the app in different contexts, such as when multitasking or when used by family members, allows developers to tailor the user experience to effectively accommodate these varied needs. For example, some participants noted that they could use the app together with their family members (FP 8,9,10, IP 1,3,4,5,7). In these cases, users have to ”look over the health data”, ”understand the features of the app”, and ”set up some reminders in the app” for their family members. This additional layer of complexity emphasises the multifaceted need for generated adaptations. IP7, for example, expressed the desire to understand why specific adaptations are made for their parents, which might not align with the parents’ priorities of using the app. Similarly, IP5, FP8, and FP10 mentioned the need to monitor their parents’ health while their parents are in a different country. This usage pattern in which the app is used by individuals other than the intended user, can potentially lead to unforeseen consequences (Lewis and Wyatt, 2014), such as misinterpreting data, incorrect health decisions, and feature misuse, especially when the primary user’s health needs differ from those of the secondary user (Aljedaani et al., 2021; Lewis and Wyatt, 2014). This interaction between users and potential users, including informal carers (e.g. children, spouses) and formal carers (e.g. doctors, nurses), is also highlighted in the study by Martin-Hammond et al. (2019).
4.3. Strategies to mitigate AUI challenges and relevant moderators
Participants discussed strategies relating to controllability, user support, and alignment need for designing adaptations during our user study session (summarised in Table 3). Participants exhibited varied priorities for different strategies, influenced by their preferred level of involvement in the system, their experience with mHealth technologies, and their own health conditions (Moderators).
| Categories | Concepts | Participants | #(%) | |
| Strategies | S1. Controllability and Autonomy (9 FP, 7 IP) | Dedicated adaptation panel to manage all adaptations | FP 7, IP 3,7 | 3/22 (14%) |
| Dedicating a specific part of the interface for adaptation | FP 4,6,9, IP 1,4,5 | 6/22 (27%) | ||
| Step-by-step adaptation process | FP 3,10,11,12, IP 1,3 | 6/22 (27%) | ||
| Configuring factors for different adaptations | FP 3,7,9, IP 2,3,4-7 | 9/22 (41%) | ||
| Accessible opt-in and opt-out process for the adaptation | FP 3,5,7,11,12,15, IP 1,3,4,6,7 | 11/22 (50%) | ||
| S2. User support for adaptation (6 FP, 6 IP) | Shortcut to different adaptations | FP 10,11,13, IP 2,3 | 5/22 (23%) | |
| Onboarding processes, educational materials & tutorials | FP 1,7,9, IP 1,5,7 | 6/22 (27%) | ||
| Provide user resource to personalise their experience | FP 7,9, IP 3,4,7 | 5/22 (23%) | ||
| Contextual explanations of adaptations | FP 7,8,9,13, IP 3,4 | 6/22 (27%) | ||
| S3. Alignment (6 FP, 10 IP) | Alignment of user’s chronic diseases | FP 1,2,3,4,6,9, IP 1,3 | 8/22 (36%) | |
| Alignment of user’s motivation and goal | FP 5,8,13, IP 3,4,7 | 6/22 (27%) | ||
| Alignment of existing technology and guideline | FP 3,4,6,7, IP 1,4,6 | 7/22 (32%) | ||
| Moderators | M1. User involvement | Active/passive user involvement | All participants | 22 (100%) |
| User-caregiver collaboration | FP 8,9,10, IP 1,3,4,5,7 | 8/22(36%) | ||
| M2. User experience with mHealth applications | All participants | 22 (100%) | ||
| M3. Health condition | All participants | 22 (100%) | ||
4.3.1. S1. Controllability and Autonomy
Participants in the study stressed the importance of having control and autonomy over the adaptation process. While autonomy and controllability are distinct concepts, they are closely related and can influence each other in the context of user interactions (Jameson, 2007). Some participants expect to have a dedicated adaptation panel that allows them to manage all adaptations (FP 7, IP 3,7). In this panel, they envision having comprehensive control over various aspects of adaptation, enabling them to fine-tune their user experience. For instance, IP3 explained:
In contrast to controlling adaptation in multiple places on the UI, some participants preferred to have a specific part of the system dedicated to adaptation which introduces only minimal distortion to the original interface (FP 4,6,9, IP 1,4,5). This preference resonates with the insights gained from previous studies (Gajos et al., 2006; Li and Gajos, 2014; Deuschel and Scully, 2016), possibly arising from the inherent desire to maintain high levels of spatial stability in the UI (Deuschel, 2018). By centralising adaptation controls in a specific part of the system, users can efficiently manage and customise adaptations without compromising the integrity of the overall interface (Deuschel, 2018). IP1 explained that she would feel more confident exploring the system’s adaptation features if there was a designated adaptation section within the app, as it would prevent ”accidentally making changes to different parts of the app”. This concern about unintentional changes is not uncommon among some elderly users (mentioned by FP 4,6,9), who may feel anxious about unintentionally altering the adaptation settings, often requiring assistance to address such issues. The confidence older individuals have in their technological proficiency significantly influences their adoption of mHealth (Cajita et al., 2018; Bentley et al., 2014).
One challenge previously highlighted is that users may not initially know which adaptation suits them best. Participants identified a potential solution by providing a step-by-step adaptation process that gradually introduces more complex adaptations (IP 1,3, FP 3,10,11,12). This approach aligns with research indicating that people tend to choose tasks of intermediate difficulty, which can improve their sense of competence (Deci and Ryan, 2013). However, it is essential to adapt in an unobtrusive manner that remains easy for users to notice and follow, as some participants expressed concerns regarding the engagement of older individuals (aged over 65) with frequent adaptation (FP 4,6). These different kinds of controllability and autonomy are achieved on various levels of granularity which vary from broad UI presentation changes to more fine-grained adaptations that impact dialogues, navigational flows, or specific content elements (Abrahão et al., 2021; Cristea and Calvi, 2003; Findlater and McGrenere, 2010). Our study uncovered a spectrum of preferences for the level of adaptation granularity. Some favoured one-time adaptations that encompassed the entire application, while others wanted specific adaptations targeting particular interaction objects.
The participants emphasised the need for flexibility in configuring multiple factors for different adaptations (IP 2,3,4-7, FP 3,7,9). Offering ”user control over the automation level and scope of adaptations” (IP 3), ”capability to prioritise UI elements based on their frequency, significance or sequence of usage” (IP 2), and ”options to mark specific adaptations as favourites for increased use or for similar adaptation scenarios” (FP 9) are noted by participants. Providing users with an easy and accessible opt-in and opt-out process for adaptation is crucial to ensure user autonomy and positive user experience (IP 1,3,4,6,7, FP 3,5,7,11,12,15). Participants desire a system that allows them to easily enable adaptations with a simple click (FP 3,7,11, IP 4), while also valuing the ability to revert adaptations if needed (FP 3,5). These practices for enhancing controllability empower users to maintain control over their app experience, they are more likely to remain engaged with the application over time, leading to greater user satisfaction and improved overall outcomes (O’Brien and Toms, 2008).
4.3.2. S2. Providing user support for adaptation
It is essential to provide comprehensive user support and resources to help navigate and understand adaptations. We identified two phases for this support: prior to adaptation and during adaptation. In the phase, before adaptation, users should have access to clear and detailed information about available adaptations, their potential benefits, and how to enable or disable them. Participants expressed the need for the system to offer straightforward and easily accessible shortcuts for making these adaptations, avoiding navigating through complex menus (IP 2,3, FP 10,11,13). For example, IP2 shared his experience of not noticing the adaptation feature in a mental health app for a year because it is buried deep within the settings menu. Additionally, incorporating onboarding processes or interactive tutorials can effectively introduce users to the concept of adaptation, demonstrate its benefits, and facilitate a better understanding of the underlying principles and purposes of adaptations (FP 1,7,9, IP 1,5,7).
During the adaptation process, users may need assistance or guidance to ensure smooth implementation. The availability of ample resources and meaningful suggestions plays a significant role in the desirability of adaptations. Participants have shown their appreciation for the option of selecting from a catalogue of adaptations offered by the application (IP 3,4,7, FP 7,9). As IP7 puts it:
This approach ensures that users have access to relevant suggestions while still maintaining the flexibility to adapt the technology to their preferences. Furthermore, providing contextual explanations for adaptations within the application to help users understand the adaptation process and, interpret perceived information is essential for transparency in the adaptation process (Höök, 2000). Participants frequently expressed the need for a detailed justification regarding the relationship between user data and generated adaptations, as well as the reasons behind why the system believes that a particular adaptation is needed or desired (IP 3,4, FP 7,8,9,13). Some participants also anticipated features such as ”highlighting the adapted parts on the interface” (IP 3) or ”using animated transitions to visualise the adaptation process” (FP 8). These thoughts on adaptation have been explored in prior research (Findlater and McGrenere, 2007) and reflect the importance of transparency in the operation of mHealth applications (Glass et al., 2008; Martin-Hammond et al., 2019). For instance, FP13 emphasised the importance of being able to ”verify the actions and outcomes for adaptation” as a primary factor driving his usage of the system. However, it is important to acknowledge that not all participants share this perspective. IP2, for instance, believed that contextual reasoning for adaptations is unnecessary if users have no control over what is suggested. This finding resonates with the conclusions drawn from the study by Bunt et al. (2012), which revealed that users exhibited minimal interest in explanations when they lacked direct control over an agent’s reasoning process. Furthermore, IP7 and FP4 cautioned against overly complex reasoning, which they believed could ”confuse and distract users from their primary tasks”. This divergence in opinion highlights the need to tailor support and explanations to users’ interests and cognitive abilities (Glass et al., 2008).
4.3.3. S3. Alignment
Achieving alignment in chronic disease management involves standardising and simplifying medical terminology and self-management processes. Older participants (over 55 years old) expressed the need for the system to provide a standardised, non-academic terminology for drug names (FP 1,2,3,4,6,9, IP 1,3). This request arises from the challenge of remembering the names of medications, particularly when managing multiple chronic conditions with frequently changing prescriptions. For example, FP9, who deals with heart disease, type 2 diabetes, epilepsy, and asthma, takes several medications concurrently. He faced challenges with the inconsistency in the appearance and names of medicine. FP1, FP3, and FP7 expressed similar feelings, asking for a simple and standardised naming system to aid memory recall. Additionally, some participants expressed a desire to acquire knowledge of medical terminology to facilitate more informed conversations with healthcare providers (IP 3,7). This indicates the need for a balanced approach that simplifies terminology for better comprehension while also offering avenues for users to broaden their medical understanding. The introduction of AUIs might inadvertently result in users acquiring less domain-specific knowledge compared to nonadaptive systems (Lanier, 1995). Thus, it becomes crucial to assess users’ readiness to compromise on knowledge acquisition in favour of ease of use when formulating AUI designs. Such an approach ensures that users can navigate complex medical information effectively, while also encouraging continuous learning and engagement with the application. Furthermore, many participants (FP 1,4,6,7,9, IP 1,3) recommended the incorporation of relevant visuals related to the app’s topic to enhance user understanding and engagement. For example, IP1 proposed incorporating a doctor’s image to signify critical health issues, suggesting the potential effectiveness of visual cues in conveying important health information. However, it is essential to recognise that the introduction of anthropomorphic visuals, such as the doctor’s image, may inadvertently shift users’ perceptions of responsibility away from themselves and onto the system (Höök, 2000).
[…]

### v2

@@ 3.3.3. Quantitative data analysis
In stage two, as the surveys targeting end-users and software practitioners primarily comprised open-ended questions, the analysis predominantly relied on descriptive statistics to summarize the responses. For the user survey in stage one, quantitative data were analyzed using R444https://www.r-project.org. Descriptive statistics were employed to summarize respondent characteristics and preferences, providing insights into distribution patterns and emerging trends within the dataset. Chi-square tests were used to determine whether there was a significant association between the preferences of users with respect to various aspects of adaptations and their demographic characteristics, such as age, gender, nationality, education, and chronic diseases. To ensure that the data distribution meets the prerequisites for the Chi-square independence test, related variables were grouped into categorical variables beforehand. Age was categorized into two groups: 18-45 and 45-74. The level of education was classified into three groups: Less than Bachelor’s degree, Bachelor’s degree, and Postgraduate (Master’s and Doctoral degrees). Chronic disease conditions were categorized as detailed in Section 5.1. If a significant association is found, binary logistic regression or multinomial logistic regression will be subsequently employed to model the relationship between these variables. Understanding how these demographic factors influence user preferences provides valuable information on how to tailor AUIs to accommodate various user needs and preferences. The common significance level of is chosen for statistical analysis.
4. Stage One: Findings of interview and focus group studies
This section presents key findings on the user’s perspectives toward the AUI prototype in the context of mHealth applications. We identified four overarching challenges that participants faced while interacting with the AUI prototype. As participants describe the challenges they encountered, they also offer recommendations for improving the adaptation design. For more comprehensive insights into each challenge and recommendation, refer to Wang et al. (2024). In addition, a further detailed analysis explores why certain recommendations are more effective for some users than others, identifying three contextual factors that influence how individual user characteristics shape adaptation preferences and outcomes (Section 4.1). Drawing on STGT data analysis and the existing literature (Wang et al., 2023b; Abrahão et al., 2021), we grouped the identified challenges into four distinct categories: What to adapt, which focuses on pinpointing specific UI components that require adaptation and recognizing the associated implementation challenges; Who should initiate adaptations, which explores the assignment of responsibilities between users and systems in triggering adaptations; How to adapt, which examines the strategies and mechanisms used to carry out adaptations effectively; and When to adapt, which considers the appropriate timing and contextual conditions for initiating adaptations. Participants participated in discussions about recommendations related to controllability, user support, and alignment of adaptation design. The priorities of the participants placed on various recommendations are diverse and were shaped by their desired degree of involvement with the system, their familiarity with mHealth applications, as well as their personal health conditions (contextual factors).
4.1. Contextual Factors of Recommendations
4.1.1. User involvement
The activeness level of users significantly influences recommendations related to user participation, such as user support and controllability. In many human-computer interaction models, users are typically categorized as ”active process operators” or ”passive process operators” (Persson et al., 2001). This distinction is common in shared technologies such as health technologies used by physicians, nurses, and patients (Xu et al., 2014). Users can be active, having direct control, or passive, interacting without control (Montague and Xu, 2012; Inbar and Tractinsky, 2009). Our research categorizes users as active or passive based on their willingness to engage with the system, regardless of the shared technology (Eiband et al., 2021; Gómez-Carmona et al., 2024). Active user involvement is exemplified by participants who proactively experiment with various data sources to understand how they impact the system’s output, desire for active participation, and a sense of control over the adaptation process. They are active in the adaptation process, willing to explore and approve various adaptation suggestions (Gajos and Chauncey, 2017). On the contrary, passive user involvement characterizes individuals who are more inclined to seek information about how the adaptation works but do not actively provide feedback or corrections to the system. Passive participants may stop exploring once they believe the current UI meets their minimum requirements (Eiband et al., 2021; Parra and Brusilovsky, 2015). Participants with a passive involvement with the application, with some admitting that they have not fully explored what the application can do, others expressing a lack of concern about the system’s adaptation process, some indicating tolerance for most generated adaptations, and some preferring minimal interaction with the application. Participants can actively experiment with the way that different adaptation settings influence system output if the software provides switches and configurable options within the UI. For example, they can access a dashboard for the adaptation types and configure the settings for different adaptations. In contrast, users with a negative perception of the software show less enthusiasm for this level of participation. They tend to prefer less or even no interaction when it comes to experimenting with adaptation settings. Participants also highlighted the importance of including family members or caregivers in the use of the system, as caregivers often play an active role in medical care decisions (Xie, 2009; Cassileth et al., 1980). Allowing others to handle the adaptation process can alleviate cognitive effort for older users (Schiavo et al., 2022), but could potentially reduce their independent use of the technology (Munteanu and Salah, 2017).
4.1.2. User experience with mHealth applications
[…]